### v1

@@ Abstract
Computational notebooks have become the primary coding environment for data scientists. Despite their popularity, research on the code quality of these notebooks is still in its infancy, and the code shared in these notebooks is often of poor quality. Considering the importance of maintenance and reusability, it is crucial to pay attention to the understandability of the notebook code and identify the notebook metrics that play a significant role in their understandability. The level of code understandability is a qualitative variable closely associated with the user’s opinion about the code. Traditional approaches to measuring it either use limited questionnaires to review a few code pieces or rely on metadata such as likes and votes in software repositories. In our approach, we enhanced the measurement of the understandability level of Jupyter notebooks by leveraging user comments related to code understandability within a software repository. As a case study, we started with 542,051 Kaggle Jupyter notebooks, named DistilKaggle, which were introduced in our previous research. To identify user comments associated with code understandability, we utilized a fine-tuned DistilBERT transformer. We established a user opinion-based criterion for measuring code understandability by considering the number of code understandability-related comments, the upvotes on those comments, the total views of the notebook, and the total upvotes received by the notebook. We refer to this criterion as User Opinion Code Understandability (UOCU), which has proven much more effective than previous approaches. In addition, we were able to train machine learning models which could predict code understandability of notebooks solely based on their metrics. We collected a total of 34 metrics for 132,723 final notebooks as features in our dataset, while UOCU is considered as the label for them. Using the Random Forest classifier, our predictive model achieved 89% accuracy in predicting code understandability levels in computational notebooks.
@@ 1 Introduction
Computational notebooks have become the primary coding environment for data scientists psallidas2019data . They offer several advantages over traditional software development environments, including easy documentation, code sharing, result analysis, visual and intuitive code development, and the ability to compile, execute, and modify code cells without re-executing the entire notebook. The advantages outlined above establish computational notebooks as a fundamental tool in the development of data analysis solutions. For instance, a case study at Microsoft highlighted that data scientists rely on computational notebooks for sharing and reusing their code across software teams epperson2022strategies . In another empirical study, cleaned computational notebooks were introduced as a best practice for a friction-less transition of data-driven projects from the exploration stage to production as a software engineering product quaranta2022assessing .
@@
The quality and understandability of implementations in computational notebooks are crucial for various purposes, such as education to provide coding best practices for future data scientists wang2020better , notebook reusability epperson2022strategies , and maintainability wang2020assessing ; yang2021subtle ; dong2021splitting . Recent studies have focused on improving the quality of computational notebooks through enhanced documentation liu2019understanding ; rule2018exploration ; wang2022documentation ; yang2021subtle ; venkatesh2023enhancing , notebook structure wenskovitch2019albireo ; titov2022resplit ; jiang2022elevating , and managing notebook variants and revisions subramanian2019supporting ; jiang2022elevating . However, efforts to enhance code understandability in computational notebooks are still in the early stages venkatesh2021automated . This paper aims to address this problem.
@@
Code Understandability (CU), also referred to as program understandability or code comprehensibility, is the active process by which developers acquire knowledge about a software system through exploring software artifacts, reading source code, and consulting documentation xia2017measuring . On average, developers spend approximately 60% of their time engaged in program comprehension during software development and maintenance xia2017measuring . To measure CU, various approaches have been proposed, often involving gathering feedback from individuals regarding a program or specific code segments. These studies employ different response variables, including time, accuracy, opinion, and visual metrics, as criteria for assessing CU. Some studies directly solicit opinions from a limited number of developers about specific code pieces sykes1983effect ; buse2009learning ; medeiros2018investigating ; scalabrino2019automatically . Other studies rely on metrics such as likes, stars, or votes on software repositories to gauge CU lu2018internal ; nasehi2012makes ; wang2022documentation ; liu2021haconvgnn . However, findings from small-scale studies may not be widely applicable or provide a comprehensive understanding of how developers perceive code. Additionally, user votes on software projects may be influenced by factors beyond CU, such as ease of use, algorithm accuracy, popularity, and marketing considerations.
@@
In this paper, we have employed the opinion response variable oliveira2020evaluating to enhance the measurement of CU in a unique manner. Instead of conducting surveys, direct inquiries, or relying solely on user votes, we have developed a novel approach that involves extracting and analyzing comments regarding the understandability of notebooks in a repository. As a case study, we chose Kaggle [^1] https://www.kaggle.com, which is a notebook repository with a variety of interactive and collaborative functionalities. In this regard, we began by obtaining the metadata files of the Jupyter notebooks and then applied preprocessing to the user comments. Using a fine-tuned DistilBERT transformer, we identified user comments related to CU and provided a metric considering how much each comment attracted other users’ attention. We named this metric User Opinion Code Understandability (UOCU). Compared to a dataset labeled by 42 experts, UOCU has been shown to outperform current state-of-the-art methods for measuring the understandability of notebooks, as demonstrated in our evaluation. Incorporating the criterion of notebook upvotes in UOCU metric yields demonstrably superior outcomes in our evaluations, as evidenced by our research findings.
@@
Another contribution of our study is the exploration of notebook metrics that have a more significant impact on CU. Previous studies have introduced a set of code metrics (such as lines of code, maximum line length, and cyclomatic complexity) that are related to CU for traditional programming languages like Java, Python, and C buse2009learning ; posnett2011simpler ; kasto2013measuring ; scalabrino2019automatically ; grotov2022large . However, some of these metrics (such as coupling, cohesion, and depth of inheritance tree) are not relevant to notebooks. Notebooks may have more metrics due to structural differences compared to code scripts addressed in recent studies grotov2022large ; wang2020assessing ; wang2020better ; titov2022resplit ; ono2021interactive ; wang2021restoring ; zhu2021restoring ; dilhara2021understanding ; titov2022resplit ; yang2021subtle ; wang2022documentation ; venkatesh2023enhancing . Additionally, studies that propose some notebook metrics as effective measures for assessing CU lack a systematic approach supported by significant big data.
To address this gap, we conducted a comprehensive study of Jupyter notebook metrics to identify 34 metrics related to CU. We employed a machine learning approach to investigate the correlation between CU and Jupyter notebook metrics. For this purpose, we leveraged the DistilKaggle dataset, which comprised 542,051 notebooks shared on the Kaggle platform and was originally introduced in mostafavi2024DistilKaggle . We preprocessed KGTorrent data and distilled a dataset, extracting 34 metrics from each notebook. Finally, we merged this dataset with the output pertaining to UOCU, allowing us to analyze the relationship between the metrics and CU in a comprehensive manner.
@@
After assembling a dataset containing notebook metrics and their associated CU scores, we developed a machine learning model using four different algorithms, including CatBoost, XGBoost, Decision Tree, and Random Forest. The best result we achieved for predicting the understandability of notebook codes was with the Random Forest algorithm, which achieved an accuracy of 89% and an F1-score of 88%, improving previous criteria. With this acceptable level of accuracy and F1-score, we identified the most influential notebook metrics for CU. These metrics included developer expertise level, number of blank lines of code, nubmer of visualization data types, number of executed cells, number of H1 tags in markdown headlines and number of lines of markdown cells.
@@
  • creating a dataset for CU through fine-grained elicitation of understandability-related metrics
@@
  • generating a dataset and creating a model to predict the understandability levels within Jupyter notebooks
@@
This paper is organized as follows: Section 2 presents a motivating example. Section 3 provides a systematic review of notebook metrics. Section 4 identifies the problems and research questions in detail. Our proposed approach is described in Section 5, while Section 6 presents our evaluations and results. Section 7 discusses related work. The limitations and validity risks associated with our findings are explained in Section 8. Finally, Section 9 concludes the paper and provides future research directions.
Our dataset and experimental source code are available online [^2] https://github.com/ISE-Research/NotebookCU. Moreover, our service, which accepts a notebook and returns its code metrics and predicts its CU level is accessible [^3] http://46.249.101.150/docs.
@@ 2 Motivating Example
Some studies leverage the quantity of user votes received by a code or notebook within a software repository as an indicator of its understandability level nasehi2012makes ; lu2018internal ; wang2022documentation ; liu2021haconvgnn . However, user votes can be influenced by various factors such as output accuracy, popularity, and ease of reuse, among others. We have selected 10 notebook samples from the Kaggle repository along with users’ upvote data up to August 10, 2023 (latest upvote in Table 1). Three programming experts with at least three years of experience working with Jupyter notebooks were asked to evaluate each example and place them into either a normal or good CU group. Table 1 illustrates the average outcomes of the experts’ evaluations, representing the prevailing expert assessment of CU across the notebook samples. As can be seen, in these notebooks, the experts’ opinions contradicted the results of the user votes. Upon an initial review of the comments related to these notebooks, it is evident that the comments for the first four notebooks have little relevance to CU. Conversely, the comments regarding CU for the last four notebooks (e.g.,“nicely organized file and good EDA”, “well-explained notebook”, “very well presented and nice kernel”) are in a significantly better state, based on the number of visitors and upvotes received. Therefore, compared to the opinions of experts, these comments are considered to reflect a higher level of understandability satisfaction.
@@
Notebooks* Latest Upvote Expert Score
A 66 normal
B 72 normal
C 96 normal
D 113 normal
E 5 good
F 21 good
G 21 good
H 40 good
@@
Based on the expert opinions in Table 1 and the metrics provided in Table 2, we can hastily conclude that there is a direct relationship between some notebook metrics and CU. Specifically, sufficient documentation using markdowns and headlines, preserving the output of each code cell, and presenting results using visualization tools all contribute to a better understanding of notebook code. In this paper we aim to explore the applicability of these findings to other cases and examine additional metrics in order to assess their generalizability.
@@ 3 Metrics of Notebooks
Metrics play a vital role in evaluating the effectiveness, comprehensibility, and quality of computational notebooks. Metrics provide a systematic framework for measuring and analyzing various aspects of notebooks, enabling a thorough assessment. In this section, we delve into two distinct groups of metrics: Notebook-Based Metrics and Script-Based Metrics. Notebook-Based Metrics focus on the properties and structure of the notebook itself, including organization, documentation, and visual representation. On the other hand, Script-Based Metrics concentrate on the underlying code, examining factors such as complexity, modularity, and adherence to coding standards. Through the exploration and analysis of these metrics, our aim is to gain deeper insights into the strengths, weaknesses, and overall effectiveness of computational notebooks.
@@ 3.1 Notebook-Based Metrics
Our first set of metrics focuses on assessing the quality of Jupyter notebooks, a topic that has received considerable attention in recent years. Numerous articles have explored the evaluation of code quality in this environment, introducing a variety of metrics grotov2022large ; wang2020assessing ; wang2020better ; titov2022resplit ; ono2021interactive ; wang2021restoring ; zhu2021restoring ; dilhara2021understanding ; yang2021subtle ; wang2022documentation ; venkatesh2023enhancing ; liu2021haconvgnn . We conducted an extensive review of literature from top-tier journals and conferences in software engineering and data science domains, employing snowballing techniques to augment our search and uncover additional relevant contributions. In selecting metrics related to CU (CU), our approach was guided by the comprehensive definition provided by Xia et al. (2017), emphasizing the active process through which developers acquire knowledge about software systems, particularly within the Jupyter notebook environment. This definition underscores activities such as exploring software artifacts, reading source code, and consulting documentation. Leveraging insights from domain experts with extensive expertise in Python development, averaging 5 years of experience in software engineering or data science, we refined an initial list of 23 metrics down to 17, prioritizing those most pertinent to CU in Jupyter notebooks. These metrics are detailed in Table 3. Subsequently, we will elaborate on metrics whose names may not clearly indicate their conceptual underpinnings.
@@
[b] Feature Abbreviation Reference Script-Based Metrics Basic Metrics Lines of Code LOC  scalabrino2019automatically ; buse2009learning ; kasto2013measuring ; wang2022documentation Number of Blank Lines of Code BLC  scalabrino2019automatically ; buse2009learning ; kasto2013measuring Lines of Comments LOCom  scalabrino2019automatically ; buse2009learning ; kasto2013measuring Number of Comment Words CW  kasto2013measuring Number of Statements S  kasto2013measuring ; scalabrino2019automatically Number of Parameters P  kasto2013measuring ; scalabrino2019automatically Number of User-Defined Functions UDF  grotov2022large Complexity Metrics Cyclomatic Complexity CyC  scalabrino2019automatically ; buse2009learning ; kasto2013measuring ; grotov2022large Nested Block Depth NBD  kasto2013measuring ; scalabrino2019automatically Kind of Line of Code Identifier Density KLCID  klemola2003cognitive Halstead Metrics Number of Operands OPRND  posnett2011simpler ; kasto2013measuring ; scalabrino2019automatically Number of Operators OPRAT  posnett2011simpler ; kasto2013measuring ; buse2009learning Number of Unique Operands UOPRND  posnett2011simpler ; kasto2013measuring Number of Unique Operators UOPRAT  posnett2011simpler ; kasto2013measuring Readability Metrics Average Line Length of Code ALLC  buse2009learning Number of Identifiers ID  buse2009learning Average Number of Identifiers AID  buse2009learning Notebook-Based Metrics Number of Code Cells CC  pimentel2019large ; wang2022documentation Mean Number of Lines in Code Cells MeanLCC  kasto2013measuring Number of Imports I  pimentel2019large ; kasto2013measuring Mean Number of Words in Markdown Cells MeanWMC  wang2022documentation Number of H1 tags in the Markdown H1  pimentel2019large ; venkatesh2021automated ; venkatesh2023enhancing Number of H2 tags in the Markdown H2  pimentel2019large ; wang2020assessing ; venkatesh2023enhancing Number of H3 tags in the Markdown H3  pimentel2019large ; wang2020assessing ; venkatesh2023enhancing Number of Markdown Cells MC  wang2022documentation ; pimentel2019large ; liu2021haconvgnn Mean Number of Lines in Markdown Cells MeanLMC  grotov2022large Number of Markdown Words MW  kasto2013measuring ; yang2021subtle Number of Lines in Markdown Cells LMC  scalabrino2019automatically ; wang2022documentation Distribution of Markdown Cells DMC  pimentel2019large ; wang2022documentation Distribution of Imports DI  pimentel2019large ; wang2022documentation Performance Tier PT  alomar2021behind ; scalabrino2019automatically External API Popularity EAP  scalabrino2019automatically ; zhu2021restoring ; dilhara2021understanding Number of Visualization Data Type VDT  ono2021interactive ; agrawal2022understanding Number of Executed Cells EC  pimentel2019large
Table 3: Features Extracted in this Study
@@
  • Number of Visualization Data Type: Considering that it is possible to display the output of each cell in notebooks by executing it, this metric shows the number of visual outputs, including all kinds of images and graphs.
@@
  • Number of Executed Cells: This attribute specifies the number of cells that have been executed in a notebook and have an execution order number.
  • H1, H2 and H3 headlines: These metrics show the number of headlines that are written inside the markdowns and are usually used to express titles and subtitles.
  • Distribution of Markdown Cells: For calculating the distribution of markdowns throughout a notebook, we suggest the inverse coefficient of variation. The coefficient of variation (CV), also known as relative standard deviation (RSD), is a statistical measure commonly used for comparing diversity in workgroups and is defined as the ratio of the standard deviation to the mean bedeian2000use .
    $CV=\sigma/\mu$ (1)
    where:
    $\sigma=\sqrt{\sum(x_{i}-\mu)^{2}/n}$ (2)
    where:
    - n: the size of population
    - $x_{i}$: each value from the population
    To calculate the distribution of markdown cells, we considered the number of words of each markdown as the size of it. After applying the inverse of Equation (1) to the notebooks of the final dataset, the highest CV (the most unbalanced distribution) assigned to notebook P [^4] P:https://www.kaggle.com/code/xiwuhan/jmtc-2nd-place-solution and the lowest CV (the most balanced distribution) was for notebook Q [^5] Q:https://www.kaggle.com/code/anokas/two-sigma-time-travel-eda.
  • Distribution of Imports: To calculate the amount of distribution, we used Equation (1), too. Here, we consider the size of each code cell as the number of imports in it. In this way, the highest CV was assigned to notebook R [^6] R:https://www.kaggle.com/code/moazmagdy/titanic-survival-prediction, in which, despite the high number of code cells, all import commands were accumulated in one cell and the lowest CV was assigned to S [^7] S:https://www.kaggle.com/code/arturkiulian/coronawhy-org-team-980-people-join-slack as it only contained one import command in each code cell.
@@
  • Performance Tier: This feature is present in the metadata of the Kaggle repository, which, based on specific rules, determines the level of expertise of notebook developers with one of the levels 0 to 4888https://www.kaggle.com/progression. Details of how Kaggle calculates the performance tier are given in Appendix B.
@@
  • External API Popularity: This metric assigns a number to each notebook that shows to what extent the APIs and libraries used in that notebook are popular zhu2021restoring . The higher its value, the more frequently other developers utilize the APIs featured in the notebook, which results in better CU based on previous studies scalabrino2019automatically . In order to measure this criterion in notebooks: First, we counted the number of times each API was imported in our dataset and assigned a popularity score to each API based on the frequency at which they were used in the whole dataset. Then, for each notebook, we summed up the popularity score of APIs used in that notebook to obtain the External API popularity score for that notebook.

### v2

@@ Abstract
Computational notebooks have become the primary coding environment for data scientists. Despite their popularity, research on the code quality of these notebooks is still in its infancy, and the code shared in these notebooks is often of poor quality. Considering the importance of maintenance and reusability, it is crucial to pay attention to the understandability of the notebook code and identify the notebook metrics that play a significant role in its understandability. The level of code understandability is a qualitative variable closely associated with the user’s opinion about the code. Traditional approaches to measuring it either use limited questionnaires to review a few code pieces or rely on metadata such as likes and votes in software repositories. In our approach, we enhanced the measurement of the understandability level of Jupyter notebooks by leveraging user opinions related to code understandability within a software repository. As a case study, we started with 542,051 Kaggle Jupyter notebooks, compiled in a dataset named DistilKaggle, which we introduced in our previous research. To identify user comments associated with code understandability, we utilized a fine-tuned DistilBERT transformer. We established a user-opinion-based criterion for measuring code understandability by considering the number of code understandability-related comments, the upvotes on those comments and the total views of the notebook received by the notebook. We refer to this criterion as User Opinion Code Understandability (UOCU), which has been proven to be much more effective than previous approaches. A hybrid approach combining UOCU with total upvotes further improved this criterion. Additionally, we trained machine learning models to classify notebook understandability solely based on notebook metrics. We collected 34 metrics for 132,723 final notebooks using the hybrid approach criterion. Our predictive model, built using a Random Forest classifier, achieved 89% accuracy in classifying code understandability levels in computational notebooks.
@@ 1 Introduction
Computational notebooks have become the primary coding environment for data scientists psallidas2019data . They offer several advantages over traditional software development environments, including easy documentation, code sharing, result analysis, visual and intuitive code development, and the ability to compile, execute, and modify code cells without re-executing the entire notebook. The advantages outlined above establish computational notebooks as a fundamental tool in the development of data analysis solutions. For instance, a case study at Microsoft highlighted that data scientists rely on computational notebooks for sharing and reusing their code across software teams epperson2022strategies . In another empirical study, cleaned computational notebooks were introduced as a best practice for a frictionless transition of data-driven projects from the exploration stage to production as a software engineering product quaranta2022assessing .
@@
The quality and understandability of implementations in computational notebooks are crucial for various purposes, such as education to provide coding best practices for future data scientists wang2020better , notebook reusability epperson2022strategies , and maintainability wang2020assessing ; yang2021subtle ; dong2021splitting . Recent studies have focused on improving the quality of computational notebooks through enhanced documentation liu2019understanding ; rule2018exploration ; wang2022documentation ; yang2021subtle ; venkatesh2023enhancing , notebook structure wenskovitch2019albireo ; titov2022resplit ; jiang2022elevating , and managing notebook variants and revisions subramanian2019supporting ; jiang2022elevating . However, efforts to enhance code understandability in computational notebooks are still in the early stages venkatesh2021automated . Our study aims to address this problem.
@@
Code Understandability (CU), also referred to as program understandability or code comprehensibility, is the active process by which developers acquire knowledge about a software system through exploring software artifacts, reading source code, and consulting documentation xia2017measuring . On average, developers spend approximately 60% of their time engaged in program comprehension during software development and maintenance xia2017measuring . Various approaches have been proposed to measure CU, often involving gathering feedback from individuals regarding a program or specific code segments. These studies employ different response variables, including time, accuracy, opinion, and visual metrics, as criteria for assessing CU. Some studies directly solicit opinions from a limited number of developers about specific code pieces sykes1983effect ; buse2009learning ; medeiros2018investigating ; scalabrino2019automatically . Other studies rely on metrics such as likes, stars, or votes on software repositories to gauge CU lu2018internal ; nasehi2012makes ; wang2022documentation ; liu2021haconvgnn . However, findings from small-scale studies may not be widely applicable or provide a comprehensive understanding of how developers perceive code. Additionally, user votes on software projects may be influenced by factors beyond CU, such as ease of use, algorithm accuracy, popularity, and marketing considerations.
@@
In our study, we have employed the opinion response variable oliveira2020evaluating to enhance the measurement of CU in a unique manner. Instead of conducting surveys, direct inquiries, or relying solely on user votes, we have developed a novel approach that involves extracting and analyzing comments regarding the understandability of notebooks in a repository. As a case study, we chose Kaggle [^1] https://www.kaggle.com, which is a notebook repository with a variety of interactive and collaborative functionalities. In this regard, we began by obtaining the metadata files of the Jupyter notebooks and then applied preprocessing to the user comments. We utilized DistilBERT [^2] DistilBERT is designed to be smaller, faster, and more efficient while maintaining a high level of performance in natural language processing tasks. It is particularly effective for understanding the contextual semantics of user comments, which is crucial for evaluating CU in computational notebooks. sanh2019distilbert , a distilled version of the BERT (Bidirectional Encoder Representations from Transformers) model. Using a fine-tuned DistilBERT transformer, we identified user comments related to CU and provided a metric considering how much each comment attracted other users’ attention. We named this metric User Opinion [^3] What we mean by “opinion” refers to user comments, and it has no relation to high upvotes or other metadata. Code Understandability (UOCU). Compared to a dataset labeled by 42 experts, UOCU has been shown to outperform current state-of-the-art methods for measuring the understandability of notebooks, as demonstrated in our evaluation. Our findings indicate that a hybrid metric combining UOCU and notebook upvotes is the most effective for assigning CU scores and supervising CU in machine learning models.
@@
Another key contribution of our study is the identification of notebook-specific metrics that exert a more substantial influence on CU. While prior research has established various code metrics (e.g., lines of code, maximum line length, and Cyclomatic Complexity) for evaluating CU in conventional programming languages such as Java, Python, and C buse2009learning ; posnett2011simpler ; kasto2013measuring ; scalabrino2019automatically ; grotov2022large , many of these metrics (e.g., coupling, cohesion, and depth of the inheritance tree) are not directly applicable to notebooks. Notably, less than 3.5% of notebooks incorporate classes or object-oriented programming principles, rendering the depth of inheritance tree and other class-based CU metrics ineffective for CU evaluation in Jupyter notebooks. Due to the structural differences between notebooks and traditional code scripts, additional metrics may be relevant, as highlighted in recent studies grotov2022large ; wang2020assessing ; wang2020better ; titov2022resplit ; ono2021interactive ; wang2021restoring ; zhu2021restoring ; dilhara2021understanding ; titov2022resplit ; yang2021subtle ; wang2022documentation ; venkatesh2023enhancing . Moreover, while some studies propose certain notebook metrics as promising for assessing CU, to the best of our knowledge, based on our literature review, they lack large-scale evaluation wang2022documentation , unlike our study, which evaluates CU in more than 500k notebooks. To address this gap, we identified metrics relevant to CU, building on our previous work mostafavi2024DistilKaggle , which introduced DistilKaggle—a dataset distilled from 542,051 notebooks shared on the Kaggle platform—by calculating 34 code metrics for each notebook.
@@
Using this dataset of notebook metrics and the best CU score we identified, we trained a machine learning model using four algorithms: CatBoost, XGBoost, Decision Tree, and Random Forest. The Random Forest algorithm yielded the highest accuracy for notebook CU classification, achieving an accuracy of 89% and an F1-score of 88%, improving previous criteria. With this acceptable level of accuracy and F1-score, we identified the most influential notebook metrics for CU. These metrics included developer expertise level, number of blank lines of code, number of visualization data types, number of executed cells, number of H1 tags in markdown headlines and number of lines of markdown cells. Additionally, by applying machine learning algorithms to different groups of code metrics, we concluded that notebook-specific metrics have the greatest impact on code understandability compared to more general metrics.
@@
  • presenting a ground truth and creating a model to classify CU levels within Jupyter notebooks
@@
This paper is organized as follows: Section 2 presents a motivating example, and Section 3 provides a review of notebook metrics. Section 4 details the problems and research questions addressed in this study. Our proposed approach is outlined in Section 5, and Section 6 presents our evaluations and results. Section 7 discusses the results and their potential applications, while Section 8 covers related work. The limitations and validity risks of our findings are explained in Section 9. Finally, Section 10 concludes the paper and offers directions for future research.
Our dataset and experimental source code are available online [^4] https://github.com/ISE-Research/NotebookCU.
@@ 2 Motivating Example
Some studies leverage the quantity of user votes received by a code or notebook within a software repository as an indicator of its understandability level nasehi2012makes ; lu2018internal ; wang2022documentation ; liu2021haconvgnn . However, user votes can be influenced by various factors unrelated to CU, such as output accuracy, notebook popularity, or ease of reuse.
For example, notebooks producing visually appealing results or addressing popular topics might garner more upvotes even if their code is not well-organized or documented. To explore this further, we selected 10 notebook samples from Kaggle, along with their upvote data up to August 10, 2023(latest upvote in Table 1).
Three authors of this manuscript, with expertise in Jupyter notebooks and familiarity with code comprehension research, evaluated each sample, categorizing it as either normal or good CU [^5] Following this assessment, they developed an annotation guideline to clearly define the concepts and criteria for CU classification, which supports this research and is provided in Appendix D.. The average results of their evaluations, presented in Table 1, represent their collective assessments across the notebook samples. The first four notebooks in this table, despite having higher upvotes and an average of 23,000 views, have fewer comments related to CU. In contrast, there are notebooks that, although they have high upvotes and an average of 2,500 views, contain significant comments indicating a good understanding of the notebooks by the readers. These notebooks were also labeled as having good CU by our annotators. Here are some of the comments: “Nicely organized file and good EDA”, “Well-explained notebook Very well presented”, “nice kernel, I enjoyed reading your work”, “great job Outstanding”, “neatly, well-explained Kaggle Notebook”, “Awesome notebook with impressive visualizations” and ”Excellent work with many useful charts”. This observation suggests that user opinions may serve as better indicators of code understandability than upvotes alone.
@@
Notebooks* Latest Upvote Latest View Expert Score
A 66 40125 normal
B 72 12009 normal
C 96 7551 normal
D 113 10650 normal
E 5 1010 good
F 21 4550 good
G 21 1514 good
H 40 3251 good
@@
Based on the expert opinions in Table 1 and the metrics provided in Table 2, we can hastily conclude that there is a direct relationship between some notebook metrics and CU.
We conducted a statistical analysis using a linear regression model to examine the relationship between individual notebook metrics and CU. The analysis considered CU (based on expert scores) as the dependent variable and various notebook metrics as independent variables. Results, detailed in Table 3, indicate that specific factors significantly influence CU. Notably, the inclusion of visual data types ( $\beta$=0.9383, p=0.0007) and the use of structured headlines ($\beta$=0.9405,p=0.0006) were identified as the most significant classifier of CU. While the number of executed cells ($\beta$=0.6440,p=0.0845) showed a relatively strong correlation.
Table 3: Linear Regression for Level of CU Classification
Variable Correlation Coefficient) p-value Standard Error
Lines of Code 0.1531 0.7074 0.0007
Lines of Comment 0.0427 0.9198 0.0056
Num. of Code Cells 0.1271 0.7565 0.0030
Num. of MD Cells 0.4532 0.2558 0.0130
Num. of Executed Cells 0.6440 0.0845 0.0093
Num. of Visual Data 0.9383 0.0007 0.0058
Num. of Headlines 0.9405 0.0006 0.0041
Performance Tier -0.2582 0.5366 0.1358
It appears that certain code metrics may play a significant role in understanding the code within notebooks. While this is an initial hypothesis based on a few straightforward examples, in this paper, we aim to investigate this effect on a larger dataset of notebooks and more extensive code metrics.
@@ 3 Metrics of Notebooks
Code metrics play a vital role in evaluating the effectiveness, comprehensibility, and quality of computational notebooks.
In our previous work mostafavi2024DistilKaggle , we introduced a dataset named DistilKaggle, which contains 542,051 Jupyter notebooks along with their code metrics. This dataset provides information from notebooks in the form of two separate dataframes: one for code cells and another for markdown cells. Additionally, the process of identifying code metrics in Jupyter notebooks and the methods used to extract this information are explained in detail. Here, we briefly describe the most important aspects related to this dataset while omitting specific details. We delve into two distinct groups of metrics: Notebook-Based Metrics and Script-Based Metrics. Notebook-Based Metrics focus on the properties and structure of the notebook itself, including organization, documentation, and visual representation. On the other hand, Script-Based Metrics concentrate on the underlying code, examining factors such as complexity, modularity, and adherence to coding standards. Through the exploration and analysis of these metrics, our aim is to gain deeper insights into the strengths, weaknesses, and overall effectiveness of computational notebooks.
@@ 3.1 Notebook-Based Metrics
These metrics are detailed in Table 4. Subsequently, we will elaborate on metrics whose names may not clearly indicate their conceptual underpinnings.
@@
[b] Feature Abbreviation Reference Script-Based Metrics Basic Metrics Lines of Code LOC  scalabrino2019automatically ; buse2009learning ; kasto2013measuring ; wang2022documentation ; lavazza2023empirical Number of Blank Lines of Code BLC  scalabrino2019automatically ; buse2009learning ; kasto2013measuring Lines of Comments LOCom  scalabrino2019automatically ; buse2009learning ; kasto2013measuring Number of Comment Words CW  kasto2013measuring Number of Statements S  kasto2013measuring ; scalabrino2019automatically Number of Parameters P  kasto2013measuring ; scalabrino2019automatically Number of User-Defined Functions UDF  grotov2022large Complexity Metrics Cyclomatic Complexity CyC  scalabrino2019automatically ; buse2009learning ; kasto2013measuring ; grotov2022large ; lavazza2023empirical Nested Block Depth NBD  kasto2013measuring ; scalabrino2019automatically Kind of Line of Code Identifier Density KLCID  klemola2003cognitive Halstead Metrics Number of Operands OPRND  posnett2011simpler ; kasto2013measuring ; scalabrino2019automatically Number of Operators OPRAT  posnett2011simpler ; kasto2013measuring ; buse2009learning Number of Unique Operands UOPRND  posnett2011simpler ; kasto2013measuring Number of Unique Operators UOPRAT  posnett2011simpler ; kasto2013measuring Readability Metrics Average Line Length of Code ALLC  buse2009learning Number of Identifiers ID  buse2009learning Average Number of Identifiers AID  buse2009learning Notebook-Based Metrics Number of Code Cells CC  pimentel2019large ; wang2022documentation Mean Number of Lines in Code Cells MeanLCC  kasto2013measuring Number of Imports I  pimentel2019large ; kasto2013measuring Mean Number of Words in Markdown Cells MeanWMC  wang2022documentation Number of H1 tags in the Markdown H1  pimentel2019large ; venkatesh2021automated ; venkatesh2023enhancing Number of H2 tags in the Markdown H2  pimentel2019large ; wang2020assessing ; venkatesh2023enhancing Number of H3 tags in the Markdown H3  pimentel2019large ; wang2020assessing ; venkatesh2023enhancing Number of Markdown Cells MC  wang2022documentation ; pimentel2019large ; liu2021haconvgnn Mean Number of Lines in Markdown Cells MeanLMC  grotov2022large Number of Markdown Words MW  kasto2013measuring ; yang2021subtle Number of Lines in Markdown Cells LMC  scalabrino2019automatically ; wang2022documentation Distribution of Markdown Cells DMC  pimentel2019large ; wang2022documentation Distribution of Imports DI  pimentel2019large ; wang2022documentation Performance Tier PT  alomar2021behind ; scalabrino2019automatically External API Popularity EAP  scalabrino2019automatically ; zhu2021restoring ; dilhara2021understanding ; mostafavi2024can Number of Visualization Data Type VDT  ono2021interactive ; agrawal2022understanding Number of Executed Cells EC  pimentel2019large
Table 4: Jupyter Notebook Code Metrics mostafavi2024DistilKaggle
@@
  • Number of Visualization Data Types: This metric reflects the total number of visual outputs displayed by executing each cell in the notebooks. It encompasses various types of visual outputs, including images, graphs, and plots.
@@
  • Number of Executed Cells: This attribute indicates the total number of cells that have been executed in a notebook, each accompanied by its respective execution order number.
  • H1, H2, and H3 Headlines: These metrics indicate the number of headings present within the markdown cells, typically used to denote titles and subtitles.
  • Distribution of Markdown Cells: This metric refers to the way these cells are spread out throughout the content, reflecting how often they appear and how substantial their content is in terms of word count. A higher degree of imbalance in this distribution suggests a concentration of markdown content in fewer cells, while a more balanced distribution indicates a more even spread of markdown content across multiple cells.
@@
  • Performance Tier: This feature is included in the metadata of the Kaggle repository and determines the level of expertise of notebook developers based on specific criteria, categorized into levels ranging from 0 to 4666https://www.kaggle.com/progression. Detailed information on how Kaggle calculates the performance tier can be found in Appendix A.
@@
  • External API Popularity: This metric assigns a score to each notebook based on the popularity of the APIs and libraries used within it zhu2021restoring . A higher score indicates that other developers frequently utilize the APIs featured in the notebook, which has been shown to correlate with improved CU in previous studies scalabrino2019automatically .